# Enhancing the De-identification of Personally Identifiable Information in Educational Data

### v1

@@ IV-F Generalizability Analysis
To further evaluate the performance of different models for PII detection, we further used the TSCC dataset (as introduced in Section III-A) to evaluate the models that were implemented on the CRAPII dataset. Table X presents the performance for different PII detection models on the TSCC dataset. The metrics are the same as those presented in Table VII.
We selected one Presidio model Presidio (en_core_web_trf) model as it demonstrated higher precision and recall compared to the en_core_web_lg variant (Model 1 from Table VII). As shown in Table X, the Presidio (en_core_web_trf) model demonstrates relatively high recall (0.9368) and low precision (0.3596). This is likely due to its conservative approach to entity detection, which enables it to capture a wide range of entities, including rare or unconventional names. This conservatism results in a large number of false positives (2,694), the highest among all six models, which significantly impacts its precision.
@@
Next, we investigated the fine-tuned GPT-4o-mini model which was fine-tuned on the CRAPII dataset. To gain a better understanding of the GPT models on PII detection on TCSS dataset. We used a three-shot prompting strategy to directly prompted GPT-4o-mini to identify PII entities in the TSCC dataset. We also prompted the Fine-tuned GPT-4o-mini model (fine-tuned on the CRAPII dataset, Model 5 in Table VII) without examples to assess its generalizability on the unseen TSCC dataset. Both models employed the same prompt structure presented in Table III, adjusting the user prompt to only label names with special identifiers as names were the only PII category in the dataset. The results in Table X show that directly prompting GPT-4o-mini with few-shot examples achieves the highest recall among all models at 0.9932, demonstrating the model’s capability to detect almost all true PII entities in the dataset. However, its precision is relatively low at 0.7145 due to the higher number of false positives (641). While this approach achieves a strong $F_{5}$ score of 0.9785, the $F_{1}$ score of 0.8311 indicates that the lower precision affects its overall performance. Then, prompting the Fine-tuned GPT-4o-mini model (on the CRAPII dataset) results in the highest precision across all models (0.9984), with only two false positives. This indicates that the entities it detects are highly accurate. However, the recall is relatively low at 0.7882, leading to a $F_{1}$ score of 0.8810 and a lower $F_{5}$ score of 0.7947. This approach is particularly suited for scenarios where precision is more critical than recall.
@@
We also provided users with a practical option to improve model performance by fine-tuning on a minimal labeled subset of their dataset, enabling the model to better align with the specific characteristics of the TSCC dataset. Of the 260 transcripts in the processed TSCC dataset, we randomly selected 10 transcripts for fine-tuning purposes introduced below and used the remaining 250 transcripts for evaluation across all models. First, we directly fine-tuned GPT-4o-mini on the 10 selected transcripts. Second, we further fine-tuned the Fine-tuned GPT-4o-mini model (Model 5 from Table VII) using the same 10 transcripts. Both fine-tuning models used the prompt structure described in Table IV. As with prompting, the only adjustment was to instruct the model to label names using @@@ and ###. Fine-tuning GPT-4o-mini on the 10 selected TSCC transcripts achieves a strong balance between precision (0.9836) and recall (0.9666). With only 26 false positives and 54 false negatives, this model achieves a $F_{1}$ score of 0.9750 and a $F_{5}$ score of 0.9672. This demonstrates the potential of fine-tuning even on a small labeled subset to adapt the model to new datasets effectively. Then, further fine-tuning of the Fine-tuned GPT-4o-mini model (on the CRAPII dataset) using the 10 selected transcripts results in the best overall performance. This model achieves a high recall of 0.9895 and a precision of 0.9708, striking an excellent balance between the two metrics. Although its recall is slightly lower (about 0.3%) than the few-shot prompting model, its precision improves significantly by approximately 26%. This leads to the highest $F_{1}$ score (0.9801) and the $F_{5}$ score (0.9887) among all models, making it the most robust approach to PII detection on the TSCC dataset.
@@
V Limitations and Future Works
While our study highlights the potential of fine-tuned GPT-4o-mini models for cost-effective and accurate PII detection in educational texts, there are opportunities for further refinement and exploration. First, our evaluation primarily addressed PII categories that are universally recognized by widely used baseline methods (e.g., Microsoft Presidio and Azure AI Language). Extending this approach to more granular or domain-specific categories, such as addresses, student ID numbers, or indirect identifiers like dates, times, and ages, could benefit from additional training data and tailored customization of LLM-based methods. Second, while we focused on Microsoft Presidio and Azure AI Language for benchmarking due to their prevalence, incorporating other models or framework may provide deeper insights into the scalability and versatility of our approach. Third, although we conducted a preliminary generalizability test on the TSCC dataset, the diverse nature of educational contexts suggests that some unseen PII distributions may still present challenges. A more comprehensive evaluation across a broader range of datasets could provide further clarity on the model’s adaptability. Fourth, while the Verifier models we developed significantly enhance precision, their trade-off in recall suggests potential for further innovation, such as multi-stage verification strategies that integrate recall-oriented detection with context-driven verification. Finally, although this study prioritized optimizing detection and verification processes, exploring advanced red-teaming or adversarial testing could further bolster model robustness and resilience against potential vulnerabilities. These extensions would help ensure the security and reliability of PII detection systems across the broad and varied landscape of educational data.
VI Conclusion
This study highlights the cost-effectiveness and strong performance of fine-tuned GPT-4o-mini for identifying personal information in educational texts. Compared to rule-based (Presidio) and cloud-based (Azure AI Language) methods, the fine-tuned GPT-4o-mini model achieves notably high recall (more than 0.95) while maintaining solid precision. This balance in performance is critical for education, where missing any PII can pose privacy risks, and unneeded removals reduce research and pedagogical utility.
Verifier models further enhance precision by verifying detected entities in context, helping users preserve semantic meaning in approaches such as Hidden in Plain Sight. Our cost analysis shows that GPT-4o-mini is at least 4 times cheaper and thus much more affordable than commonly used commercial services, lowering barriers for privacy-preserving research in education. Testing on an unseen teacher-student corpus confirms that GPT-based models generalize well, even with limited labeled data for fine-tuning.
In general, the fine-tuned GPT-4o-mini model offered the most consistent overall performance and cultural generalizability. These findings suggest that carefully refined LLMs can safeguard privacy while preserving meaningful content for learning analytics and educational research at scale.
Acknowledgment
The authors extend their sincere gratitude to Yuting Wang, Andra Liu, Owen Lalis, and Shivang Gupta, all current students and faculty at Carnegie Mellon University, for their invaluable insights and suggestions on the initial development of our study.
Funding Information
@@ Acknowledgment
This research was funded by the Richard King Mellon Foundation (Grant #10851) and the Learning Engineering Virtual Institute (https://learning-engineering-virtual-institute.org/). The opinions, findings, and conclusions expressed in this paper are those of the authors alone.

### v2

@@ 4.6 Generalizability Analysis
To further evaluate the performance of different models for PII detection, we further used the TSCC dataset (as introduced in Section 3.1) to evaluate the models that were implemented on the CRAPII dataset. Table 11 presents the performance for different PII detection models on the TSCC dataset. The metrics are the same as those presented in Table 7.
We selected one Presidio model Presidio (en_core_web_trf) model as it demonstrated higher precision and recall compared to the en_core_web_lg variant (Model 1 from Table 7). As shown in Table 11, the Presidio (en_core_web_trf) model demonstrates relatively high recall (0.9368) and low precision (0.3596). This is likely due to its conservative approach to entity detection, which enables it to capture a wide range of entities, including rare or unconventional names. This conservatism results in a large number of false positives (2,694), the highest among all six models, which significantly impacts its precision.
@@
Next, we investigated the Fine-tuned GPT-4o-mini model which was fine-tuned on the CRAPII dataset. To gain a better understanding of the GPT models on PII detection on TCSS dataset. We used a three-shot prompting strategy to directly prompted GPT-4o-mini to identify PII entities in the TSCC dataset. We also prompted the Fine-tuned GPT-4o-mini model (fine-tuned on the CRAPII dataset, Model 5 in Table 7) without examples to assess its generalizability on the unseen TSCC dataset. Both models employed the same prompt structure presented in Table 3, adjusting the user prompt to only label names with special identifiers as names were the only PII category in the dataset. The results in Table 11 show that directly prompting GPT-4o-mini with few-shot examples achieves the highest recall among all models at 0.9932, demonstrating the model’s capability to detect almost all true PII entities in the dataset. However, its precision is relatively low at 0.7145 due to the higher number of false positives (641). While this approach achieves a strong $F_{5}$ score of 0.9785, the $F_{1}$ score of 0.8311 indicates that the lower precision affects its overall performance. Then, prompting the Fine-tuned GPT-4o-mini model (on the CRAPII dataset) results in the highest precision across all models (0.9984), with only two false positives. This indicates that the entities it detects are highly accurate. However, the recall is relatively low at 0.7882, leading to a $F_{1}$ score of 0.8810 and a lower $F_{5}$ score of 0.7947. This approach is particularly suited for scenarios where precision is more critical than recall.
@@
We also provided users with a practical option to improve model performance by fine-tuning on a minimal labeled subset of their dataset, enabling the model to better align with the specific characteristics of the TSCC dataset. Of the 260 transcripts in the processed TSCC dataset, we randomly selected 10 transcripts for fine-tuning purposes introduced below and used the remaining 250 transcripts for evaluation across all models. First, we directly fine-tuned GPT-4o-mini on the 10 selected transcripts. Second, we further fine-tuned the Fine-tuned GPT-4o-mini model (Model 5 from Table 7) using the same 10 transcripts. Both fine-tuning models used the prompt structure described in Table 4. As with prompting, the only adjustment was to instruct the model to label names using @@@ and ###. Fine-tuning GPT-4o-mini on the 10 selected TSCC transcripts achieves a strong balance between precision (0.9836) and recall (0.9666). With only 26 false positives and 54 false negatives, this model achieves a $F_{1}$ score of 0.9750 and a $F_{5}$ score of 0.9672. This demonstrates the potential of fine-tuning even on a small labeled subset to adapt the model to new datasets effectively. Then, further fine-tuning of the Fine-tuned GPT-4o-mini model (on the CRAPII dataset) using the 10 selected transcripts results in the best overall performance. This model achieves a high recall of 0.9895 and a precision of 0.9708, striking an excellent balance between the two metrics. Although its recall is slightly lower (about 0.3%) than the few-shot prompting model, its precision improves significantly by approximately 26%. This leads to the highest $F_{1}$ score (0.9801) and the $F_{5}$ score (0.9887) among all models, making it the most robust approach to PII detection on the TSCC dataset.
@@
5 Discussion and Conclusion
Our study demonstrates notable performance and cost-effectiveness of a Fine-tuned GPT-4o-mini model for identifying personally identifiable information in educational texts. Compared to established rule-based (Presidio) and commercial (Azure AI Language) methods, our fine-tuned model achieves a recall of over 95%, providing a viable approach for privacy protection. To address the need for high precision in downstream applications (e.g., curating anonymized datasets to support machine learning and educational research), our study also introduced a verifier model capable of filtering false positives, thus helping to preserve the semantic meaning of data when using methods such as Hidden in Plain Sight.
Furthermore, our fairness analysis reveals that the Fine-tuned GPT-4o-mini model delivers balanced performance across cultural and gender groups, reducing the biases present in baseline systems. Testing on a unseen, different data corpus confirms the model’s ability to generalize, as it requires only limited fine-tuning to achieve similar performance. Finally, our cost analysis shows that this strong performance is achievable at a lower cost than other commercial services, reducing barriers to privacy-preserving research in education.
5.1 Considerations for Deploying Proprietary and Open-Source Models
Risk of PII leakage. While our study demonstrates notable performance of using the proprietary model, i.e., GPT-4o-mini from OpenAI, the handling of sensitive data raises relevant considerations. Since data must be transmitted to OpenAI’s servers, there is a possible risk that PII, if used in training, could be exposed through techniques like prompt injection attacks, as highlighted by recent studies [Nasr et al. (2025)]. However, it should be noted that OpenAI has implemented data processing agreements and maintains a policy ensuring that data submitted via their API is not used for model training. [^13] https://openai.com/policies/data-processing-addendum These measures provide an added layer of safeguard, reducing the risk of unintentional data leakage. In comparison, while deploying open-source models locally ensures that no third-party data processors gain access to PII, it still carries baseline privacy risks. Open-source models downloaded from platforms like Hugging Face may contain obfuscated harmful code that executes during model use, allowing attackers to perform reverse shell attacks to gain control of the user’s machine and access PII, leading to further damage [Zhao et al. (2024)].
Implementation Cost. Open-source LLMs (e.g., LLaMA 3, Mistral, DeepSeek, etc.) give developers full control over their data and fine-tuning process, making them a suitable option for privacy-sensitive tasks such as educational PII de-identification. However, adopting these models also introduces implementation challenges. A typical 8-billion-parameter model (e.g., LLaMA 3.1 8B [Grattafiori et al. (2024)]) needs more than 90 GB of GPU memory to fine-tune—beyond what a single GPU can provide. This constraint pushes teams toward multi-GPU clusters and distributed training, both of which require specialized skills in parallel computing. Although memory saving methods such as LoRA and ZeRO can reduce the requirement to roughly 6 GB of VRAM, configuring them correctly still requires technical knowledge in large-model optimization [Singh et al. (2024)]. In short, open-source models are scalable and secure once deployed, but they are best suited for teams with sufficient hardware resources and technical experience. In comparison, commercial APIs such as GPT and Claude offer easier deployment and minimal infrastructure requirements, making them more accessible for small teams or educational institutions.
Energy footprint. Beyond implementation costs, the operational energy consumption of LLMs raises environmental concerns due to the associated carbon footprint. Measuring the energy footprint is challenging for proprietary models due to limited public information on model architecture and deployment, making it difficult to assess the energy costs borne by providers. While direct measurement of energy consumption is difficult, many AI models are deployed on cloud computing platforms, which allows for indirect estimation. A study measuring the energy consumption of various models found that GPT-4o-mini is among the most efficient of OpenAI’s models, with an energy cost comparable to open-source LLaMA models [Jegham et al. (2025)]. This study noted that lower electricity usage does not always equal to superior energy efficiency: models that cost less energy to run may perform worse. This distinction matters in educational PII detection, as prior work confirms that LLaMA 3.1 8B achieves suboptimal performance, particularly in terms of precision [Zambrano et al. (2025)]. Therefore, a model like GPT-4o-mini may be a reasonable choice, balancing a manageable energy footprint with the high accuracy required for effective PII de-identification.
5.2 Implications
5.2.1 Integrating Privacy-Preserving Models in AI Educational Platforms
Our study explores potential solutions to the rising concerns about privacy and data security in the deployment of AI-powered educational technologies [Ive et al. (2024)], such as chatbots. We recommend that future developments of educational tools—particularly those integrating chatbots into teaching and learning environments—consider adopting privacy-preserving models within their system development processes. Many teachers and students may unknowingly expose sensitive personal information—such as their name, age, gender, and occupation—when interacting with these AI-driven platforms [Ive et al. (2024)]. This is especially likely when users are prompted to share such information in exchange for more personalized responses from the chatbot [Carmichael et al. (2022)], a common strategy in prompt engineering to enhance the accuracy of AI outputs. Moreover, as some educational tools (e.g., chatbots) collect data for performance enhancement or research purposes, there is a risk that this data could be further exposed or misused [Nasr et al. (2025)]. Given the increasing reliance on data for improving AI systems, there is a growing need to support safer data handling practices in educational tools. In such cases, privacy-preserving techniques—such as using a PII detection model like the one explored in this study in combination with the Hidden in Plain Sight method—can help replace detected PII with synthetic entities that preserve the semantic structure of the original content. This approach may enable broader use of AI tools, including chatbots, in educational contexts without exposing sensitive personal information.
5.2.2 Expanding the Evaluation of PII De-Identification in Educational Data Mining
Our findings suggest that relying solely on aggregate performance metrics may be insufficient for thoroughly evaluating PII de-identification systems. Such metrics can obscure underlying flaws in a model’s behavior. We observed that these limitations can have ethical and practical implications, particularly in educational settings where fairness is important. To address this, we recommend that researchers and practitioners consider incorporating additional evaluation strategies—such as (1) performance breakdowns by demographic subgroups to assess fairness (e.g., equality of odds and equal opportunity), and (2) evaluations on unseen datasets to evaluate generalizability. By adopting more comprehensive evaluation practices, the reliability and fairness of PII de-identification systems can be improved for real-world educational applications.
5.3 Future Works
First, our evaluation primarily addressed PII categories that are commonly recognized by existing baseline methods (e.g., names, emails). Extending this approach to more granular or domain-specific categories—such as street addresses, student ID numbers, or indirect identifiers like dates and ages—poses additional challenges. Identifying these categories often requires deeper contextual understanding; for example, a string of digits could represent an ID number or an unrelated quantity, and an age is only sensitive when associated with an identifiable individual. Successfully detecting such PII would require the creation of new, high-quality annotated datasets and the development of more context-aware fine-tuning strategies.
Second, while we focused on Microsoft Presidio and Azure AI Language for benchmarking due to their prevalence, a broader comparative analysis would be valuable for future work. Incorporating additional frameworks—such as Google’s Cloud Data Loss Prevention API, Amazon Comprehend, or open-source models from platforms like Hugging Face—could help further clarify the strengths and limitations of our approach. Such comparisons may also highlight performance-cost trade-offs and reveal cases where alternative models perform better for specific data types.
Third, although we conducted a preliminary generalizability test on the TSCC dataset, the diverse nature of educational contexts suggests that some unseen PII distributions may still pose challenges. A more comprehensive evaluation across a broader range of datasets, representing different age groups (K-12 vs. higher education) and context (e.g., lecture or student presentation), is needed to better define the model’s operational boundaries and identify where domain-specific fine-tuning may be necessary.
Fourth, while the verifier models we developed enhance precision, their trade-off in recall requires further improvement. The current implementation, although effective at reducing false positives, can overcorrect by eliminating true PII entities, which poses a risk in privacy-sensitive applications. Future work could explore alternative multistage verification strategies, such as implementing the verification process as a multi-model plurality or veto vote, to further reduce PII leakage. This may enable more nuanced filtering, preserving high recall while improving precision.
Fifth, although this study focused on optimizing detection and verification processes, incorporating red-teaming or adversarial testing strategies is a promising direction for future work. These methods can help uncover blind spots and edge cases that may be overlooked during standard evaluation, particularly in noisy, multilingual, or informal educational data. By simulating targeted attack scenarios or modeling real-world misuse patterns, adversarial testing can offer valuable stress tests of model assumptions and reveal subtle failure modes. Integrating such strategies into the development pipeline could improve the robustness and reliability of PII detection systems across diverse deployment contexts.
Sixth, despite their high precision, fine-tuned GPT-4o-mini models exhibit false alarms for a specific group of non-PII entities: fictional names in student essays. For example, the name “Rick” in “Personas are based on user research but tell a story about your insights. An example persona might be Rick, a 47-year-old manager struggling with his work-family-life balance” is a fictional persona from a student’s essay, yet it is incorrectly identified as PII. These errors suggest that certain fictional or example names can confuse high-precision PII detection systems. Future work could focus on better distinguishing these cases, which are especially common in educational texts. Additionally, fine-tuned GPT-4o-mini sometimes fails to detect actual PII, such as author names. For example, “Maribel Navarrete” in “Francesco 1 Maribel Navarrete Design Thinking for Business Innovation April 21” is not identified as PII. This highlights the need for more context-aware training strategies to improve the model’s handling of ambiguous or domain-specific PII.
Finally, beyond the precision–recall trade-off, our model’s robustness is limited by domain shift. When evaluated on the out-of-domain TSCC dataset, which consists of teacher–student conversations, the model showed a notable drop in recall. This performance degradation is likely due to distributional differences between TSCC and our training corpus, CRAPII, which comprises student essays. To build a more broadly applicable PII detection model, future work can prioritize fine-tuning on a larger, more diverse dataset that reflects a wider range of educational contexts. This could improve the model’s generalizability and reduce performance loss when applied to unseen domains.
Declaration of Generative AI Software Tools in the Writing Process
During the preparation of this work, the authors used ChatGPT-4o in all sections in order to polish up the language and correct any grammatical errors. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Acknowledgments
@@
This research was funded by the the Learning Engineering Virtual Institute (https://learning-engineering-virtual-institute.org/) and Richard King Mellon Foundation (Grant #10851). The opinions, findings, and conclusions expressed in this paper are those of the authors alone. The authors extend their sincere gratitude to Yuting Wang, Andra Liu, Owen Lalis, and Shivang Gupta for their invaluable insights and suggestions on the initial development of our study.